# FiDeLiS: Faithful Reasoning in Large Language Models for Knowledge Graph Question Answering

### v3

@@ Abstract
Large language models (LLMs) are often challenged by generating erroneous or hallucinated responses, especially in complex reasoning tasks. Leveraging knowledge graphs (KGs) as external knowledge sources has emerged as a viable solution. However, existing KG-enhanced methods, either retrieval-based or agent-based, encounter difficulties in accurately retrieving knowledge and efficiently traversing KGs at scale. In this paper, we propose a unified framework, FiDeLiS, designed to improve the factuality of LLM responses by anchoring answers to verifiable reasoning steps retrieved from a KG. To achieve this, we leverage step-wise beam search with a deductive scoring function, allowing the LLM to validate each reasoning step and halt the search once the question is deducible. In addition, our Path-rag module pre-selects a smaller candidate set for each beam search step, reducing computational costs by narrowing the search space. Extensive experiments show that our training-free and efficient approach outperforms strong baselines, enhancing both factuality and interpretability. Code is released at https://anonymous.4open.science/r/FiDELIS-E7FC.
@@ 1 Introduction
Large language models (LLMs) have shown impressive reasoning capabilities in tackling complex tasks Yu et al. (2024). However, the reasoning of LLMs is not always reliable and can be prone to generating outputs that are either inconsistent with real-world facts Xu et al. (2024); Huang et al. (2025) or show flawed reasoning process Li et al. (2024); Sui et al. (2024), which greatly undermine the reliability of LLMs in real-world applications Kung et al. (2023); Zhang et al. (2024).
@@
To address this issue, leveraging knowledge graphs (KGs) as external knowledge sources has emerged as a viable solution Sun et al. (2023); Ma et al. (2024); Luo et al. (2024a). Unlike traditional retrieval-augmented generation (RAG) that relies on web pages or documents Liu et al. (2024); Qian et al. (2024); Bayarri-Planas et al. (2024), KGs represent information in a structured and interconnected format, where each fact is stored as entities and relations. This format supports explicit, traceable reasoning processes Pan et al. (2023) and facilitates multi-hop reasoning through graph traversal. Moreover, each fact in a KG can be traced back to its source Sui et al. (2024); Agrawal et al. (2024), providing both context and original details, which further enhances the information authenticity and reliability of the reasoning processes.
@@
Existing KG-enhanced LLM reasoning methods face notable challenges and can be roughly categorized into two primary approaches: retrieval-based and agent-based paradigms Luo et al. (2024b). Retrieval-based methods Wang et al. (2023); Luo et al. (2024a); Baek et al. (2023) retrieve relevant KG facts to support LLM reasoning by either prompting Baek et al. (2023) or fine-tuning LLMs to learn the underlying structure of KG Luo et al. (2024a, b). These methods often suffer from incomplete or imprecise information extraction due to a lack of contextual understanding or an inability to fully capture the graph structure Luo et al. (2024b).Our error analysis of a strong retrieval-based method (i.e., Luo et al. (2024a)) in §4.3 reveals that only 67% of the generated reasoning steps are valid, with 33% containing format errors or referencing non-existent KG facts. In contrast, agent-based methods Sun et al. (2023); Ma et al. (2024) treat LLMs as interactive agents that explore KGs iteratively to construct reasoning paths and generate answers. While this approach by nature can enhance reasoning accuracy, it is computationally expensive, resulting in high latency and scalability limitations. As illustrated in Figure 1, balancing faithfulness and efficiency remains a critical challenge for existing KG-enhanced reasoning methods.
@@
To this end, we propose FiDeLiS, a unified framework designed to improve the factual accuracy and reasoning efficiency of LLMs on KGQA task. FiDeLiS anchors LLM responses to verifiable reasoning steps derived from a KG by employing two core components: (1) Deductive-Verification Beam Search (DVBS) which systematically constructs and validates reasoning paths step-by-step, ensuring logical consistency and factual correctness (discussed in §3.2). This module also prevents premature reasoning termination and incorrect path extension to ensure the validity of the generated reasoning paths. (2) Path-RAG, a retrieval-augmented mechanism that pre-selects a constrained set of candidate entities and relations for each step to mitigate computational inefficiencies. It combines semantic similarity measures with graph-based connectivity analysis to optimize the search space, significantly reducing latency without sacrificing recall or accuracy (discussed in §3.1). Extensive experiments show that our method outperform strong baselines in both accuracy and efficiency, offering a scalable, training-free solution for KG-enhanced LLM reasoning. Overall, our main contributions include:
@@
  • We propose FiDeLiS, a unified framework designed to improve the factual accuracy of LLMs by grounding reasoning paths in structured KG efficiently.
@@
  • We enable efficient, verifiable reasoning by deductively validating reasoning steps and narrowing the search space with high-quality retrieval mechanism.
@@
  • FiDeLiS performs robustly across different experiments without the need for model fine-tuning, demonstrating adaptability and scalability with improved performance on multiple benchmarks.
@@
  • By anchoring responses in verifiable reasoning paths, FiDeLiS enhances interpretability, enabling users to verify and understand each reasoning step.
@@ 2 Preliminary
To acquire valid reasoning paths, most prior studies follow the retrieval-based Li et al. (2023); Luo et al. (2024a) or agent-based Sun et al. (2023) paradigm. As indicated in Luo et al. (2024b), retrieval-based methods rely on precise additional retrievers, while agent-based methods are computationally intensive and lead to high latency. To address these issues, we propose our method, FiDeLiS, to enable both efficient and faithful reasoning over KGs.
Figure 2: An illustration of FiDeLiS. Top: The workflow of Path-RAG. An LLM first extracts key terms and generates dense embeddings that feed into the Path-RAG module. Path-RAG rapidly retrieves relevant entities and relations from a pre-embedded KG and constructs candidate reasoning steps by combining semantic similarity with graph connectivity. Bottom: The workflow of DVBS. Then, the DVBS module uses an LLM-generated planning outline to guide a beam search that builds reasoning paths step-by-step over candidates constructed by Path-RAG, with deductive verification ensuring each step logically follows the previous steps and support the user question.
@@ 3 Method
Given a complex question $q$, we first use an LLM to extract key terms from $q$ and generate dense embeddings that capture the question’s core concepts. These embeddings are input into Path-RAG module, which rapidly retrieves relevant entities and relations from a pre-embedded KG to select a smaller candidate sets for further beam search step, addressing the latency and computational burden of traditional agent-based methods. Path-RAG then constructs candidate reasoning steps by combining immediate semantic similarity with the structural connectivity of the graph, overcoming the dependence on highly precise retrievers in standard retrieval-based approaches. Next, the DVBS module employs an LLM-generated planning outline to guide a beam search that builds reasoning paths step-by-step. At each step, deductive verification checks that the accumulated reasoning steps logically supports the user question, ensuring the final reasoning path is both verifiable and accurate.
@@ 3.1 Path-RAG: Reasoning Path Retrieval-Augmented Generation
Previous agent-based methods Ma et al. (2024); Sun et al. (2023) treat LLMs as agents that iteratively interact with KGs to find reasoning paths and answer, which necessitate multiple rounds of interaction between agents and KGs and lead to high computational costs and latency. We instead propose a module, Path-RAG which iteratively pre-select a smaller candidate set to reduces the search space for exploring the potential reasoning paths from KGs. It consists of three steps and we detail the workflow as follows:
@@ 4.3 Robustness Analysis
To verify the faithfulness of our step-wise method, we conduct an error analysis regarding the whole reasoning path generation using RoG Luo et al. (2024a). We quantify the validity of reasoning path using validity ratio (VR), which is defined as the ratio of reasoning steps that existed in the KG to the total number of the reasoning steps in the output reasoning path (i.e., $\mathrm{VR}=\frac{N_{\mathrm{valid-steps}}}{N_{\text{all-steps}}}$). As shown in Figure 3, only 67% of generated reasoning steps are valid, while the remaining 33% of reasoning steps either have a format error or do not exist in the KG. This illustrates that the reasoning steps generated offer few guarantees about feasibility especially when multiple consecutive steps are combined into a reasoning path. While our method leverage step-wise verification to ensure that each of the reasoning step exist in the KG and logically connected.
Figure 3: Analysis of reasoning errors in RoG Luo et al. (2024a) over WebQSP.
@@ 5 Related Work
Large language models (LLMs) demonstrate impressive capabilities in reasoning tasks but often generate hallucinated or factually incorrect outputs, particularly in complex, multi-step scenarios Huang et al. (2025); Li et al. (2024). This unreliability reduces their effectiveness in knowledge-intensive applications. Knowledge graphs (KGs) have emerged as a solution by offering structured, verifiable data that supports transparent and multi-hop reasoning Sui et al. (2024). Unlike document-based retrieval-augmented generation approaches, KGs provide direct access to relational facts, enhancing both interpretability and traceability Chen et al. (2024).
@@
KG-enhanced reasoning methods are generally categorized into retrieval-based and agent-based models. Retrieval-based approaches, such as DeCAF Yu et al. (2023), rely on text-based retrieval to select relevant information from KGs and jointly generate answers and logical forms, but their performance can degrade without precise retrieval mechanisms. In contrast, agent-based models, like ToG Sun et al. (2023), iteratively explore reasoning paths but suffer from high computational overhead. To address these limitations, recent methods like RoG Luo et al. (2024a) and GCR Luo et al. (2024b) have sought to integrate KG structure into LLM training or decoding to improve reasoning fidelity and explanation generation. To improve the faithfulness of the LLM reasoning, KD-CoT Wang et al. (2023) verifies sub-reasoning steps through external KGs to prevent errors during inference, while NSM He et al. (2021) employs a teacher-student architecture to learn intermediate supervision signals that guide reasoning.
@@ A.1 Baselines
  • RoG Luo et al. (2024a): incorporate the underling structure of KGs into LLMs throught pre-training and fine-tuning to generate the reasoning path and explanation.
  • GCR Luo et al. (2024b) propose to integrate KG structure into the LLM decoding process to conduct graph-constrained reasoning.
@@ A.2 Datasets & Metrics
We consider three KGQA benchmark: WebQuestionSP (WebQSP) Yih et al. (2016), Complex WebQuestions (CWQ) Talmor and Berant (2018) and CR-LT-KGQA Guo et al. (2024) in this work. We follow previous work Luo et al. (2024a) to use the same training and testing splits for fair comparison over WebQSP and CWQ. The questions from both WebQSP and CWQ can be reasoned using Freebase KGs [^1] https://github.com/microsoft/FastRDFStore. To address the bias in WebQSP and CWQ, which predominantly feature popular entities and there is a likelihood that their data might have been incorporated into the pre-training corpora of LLMs, we further test our method on CR-LT-KGQA (discussed in §A.2). We use the complete dataset from CR-LT-KGQA in our experiments, as it comprises only 200 samples. Each of the question can be reasoned based on the Wikidata [^2] https://www.wikidata.org/wiki/Wikidata:Main_Page. The statistics of the datasets are given in Table 10 and Table 9. To streamline the KGs, we follow RoG Luo et al. (2024a) and utilize a subgraph of Freebase by extracting all triples that fall within the maximum reasoning hops from the question entities in WebQSP and CWQ. Similarly, we construct the corresponding sub-graphs of Wikidata for CR-LT-KGQA. We assess the performance of the methods by analyzing the F1 and Hits@1 metrics for CWQ and WebQSP, and by evaluating the accuracy for CR-LT-KGQA. The statistics of the datasets can be found in Table 9 and Table 10.
@@ A.3 Backbone LLMs & Embedding Methods
Embedding Methods. We assess the robustness of the retrieval module Path-RAG on different embedding models. We consider probabilistic ranking function like BM25333https://en.wikipedia.org/wiki/Okapi_BM25, dense retrieval using smaller language models like SentenceBERT [^4] https://sbert.net and E5555https://huggingface.co/intfloat/e5-large, and more advanced embedding model like text-embedding-3-small from OpenAI [^6] https://platform.openai.com/docs/guides/embeddings.

### v4

@@ Abstract
Large Language Models (LLMs) are often challenged by generating erroneous or hallucinated responses, especially in complex reasoning tasks. Leveraging Knowledge Graphs (KGs) as external knowledge sources has emerged as a viable solution. However, existing KG-enhanced methods, either retrieval-based or agent-based, encounter difficulties in accurately retrieving knowledge and efficiently traversing KGs at scale. In this paper, we propose a unified framework, FiDeLiS [^1] Code is released at https://github.com/Y-Sui/FiDeLiS., designed to improve the factuality of LLM responses by anchoring answers to verifiable reasoning steps retrieved from KGs. To achieve this, we leverage step-wise beam search with a deductive scoring function, allowing the LLM to validate reasoning process step by step, and halt the search once the question is deducible. In addition, we propose a Path-RAG module to pre-select a smaller candidate set for each beam search step, reducing computational costs by narrowing the search space. Extensive experiments show that our method, as a training-free framework, not only improve the performance but also enhance the factuality and interpretability across different benchmarks.
@@ 1 Introduction
Large Language Models (LLMs) have shown impressive reasoning capabilities in tackling complex tasks Yu et al. (2024); Sui et al. (2025). However, their reasoning processes are not always reliable, and can be prone to generating outputs that are either inconsistent with real-world facts Xu et al. (2024); Huang et al. (2025) or show flawed reasoning process Li et al. (2024); Sui et al. (2024). Such limitations significantly undermine the trustworthiness of LLMs in practical, real-world applications Kung et al. (2023); Zhang et al. (2024).
@@
To address this issue, leveraging knowledge graphs (KGs) as external knowledge sources has emerged as a viable solution Sun et al. (2023); Ma et al. (2024); Luo et al. (2024). Unlike traditional retrieval-augmented generation (RAG) that relies on web pages or documents Liu et al. (2024); Qian et al. (2024); Bayarri-Planas et al. (2024), KGs represent information in a structured and interconnected format, where each fact is stored as entities and relations. This format supports explicit, traceable reasoning processes Pan et al. (2023) and facilitates multi-hop reasoning through graph traversal. Moreover, each fact in a KG can be traced back to its source Sui et al. (2024); Agrawal et al. (2024), providing both context and original details, which further enhances the information authenticity and reliability of the reasoning processes.
@@
Existing KG-enhanced LLM reasoning methods face notable challenges and can be roughly categorized into two primary approaches: retrieval-based and agent-based paradigms Luo et al. (2025). Retrieval-based methods Wang et al. (2023); Luo et al. (2024); Baek et al. (2023) retrieve relevant KG facts to support LLM reasoning by either prompting Baek et al. (2023) or fine-tuning LLMs to learn the underlying structure of KG Luo et al. (2024, 2025). These methods often suffer from incomplete or imprecise information extraction due to a lack of contextual understanding or an inability to fully capture the graph structure Luo et al. (2025). Our error analysis (§4.3) of a strong retrieval-based method, RoG Luo et al. (2024) reveals that only 67% of the generated reasoning steps are valid, with 33% containing format errors or referencing non-existent KG facts. In contrast, agent-based methods Sun et al. (2023); Ma et al. (2024) treat LLMs as interactive agents that explore KGs iteratively to construct reasoning paths and generate answers. While this approach by nature can enhance reasoning accuracy, it is computationally expensive, resulting in high latency and scalability limitations. As illustrated in Figure 1, how to balance the trade-off between faithfulness and efficiency remains a critical challenge for existing KG-enhanced reasoning methods.
@@
To this end, we propose FiDeLiS, a unified framework designed to enhance both the factual accuracy and reasoning efficiency of LLMs. FiDeLiS grounds LLM responses in verifiable reasoning steps derived from a KG through two core components: (1) Deductive-Verification Beam Search (DVBS), which systematically constructs and validates reasoning paths step-by-step to ensure logical consistency and factual correctness (detailed in §3.2). This module also prevents premature termination and incorrect path extensions, thereby guaranteeing the validity of generated reasoning chains. (2) Path-RAG, a retrieval-augmented mechanism that pre-selects a constrained set of candidate entities and relations at each reasoning step to mitigate computational inefficiencies. By combining semantic similarity with graph-based connectivity analysis, Path-RAG effectively narrows the search space, significantly reducing latency without compromising recall or accuracy (discussed in §3.1). Extensive experiments demonstrate that FiDeLiS outperforms strong baselines in both accuracy and efficiency, providing a scalable, training-free solution for KG-enhanced LLM reasoning. Our main contributions are summarized as follows:
@@
  • We propose FiDeLiS, a unified framework that efficiently grounds LLM reasoning in structured knowledge graphs to improve factual accuracy.
@@
  • We enable verifiable and efficient reasoning by combining deductive validation of reasoning steps with a high-quality retrieval mechanism that constrains the search space.
@@
  • We demonstrate FiDeLiS’s robustness and scalability across multiple benchmarks without requiring any model fine-tuning.
@@
  • By anchoring responses in verifiable reasoning paths, FiDeLiS enhances interpretability and allows users to verify and understand each reasoning step from LLM reasoning.
@@ 2 Preliminary
Figure 2: An illustration of FiDeLiS. Top: The workflow of Path-RAG. An LLM first extracts key terms and generates dense embeddings that feed into the Path-RAG module. Then, Path-RAG rapidly retrieves relevant entities and relations from a pre-embedded KG and constructs candidate reasoning steps by combining semantic similarity with graph connectivity. Bottom: The workflow of DVBS. Next, the DVBS module uses LLM-generated planning to guide a beam search that builds reasoning paths step-by-step over candidates constructed by Path-RAG, with deductive verification ensuring each step logically follows the previous steps and support the user question.
@@
To acquire valid reasoning paths, most prior studies follow either retrieval-based Li et al. (2023); Luo et al. (2024, 2025) or agent-based Sun et al. (2023); Ma et al. (2024) paradigm. As indicated in Luo et al. (2025), retrieval-based methods rely on additional precise retrievers, while agent-based methods are computationally intensive and may lead to high latency. To balance this trade-off, we propose our method, FiDeLiS, to enable both efficient and faithful reasoning over KGs.
@@ 3 Method
Given a complex question $q$, we first employ an LLM to extract key terms and generate dense embeddings that capture the question’s core concepts. These embeddings serve as input to the Path-RAG module, which efficiently retrieves relevant entities and relations from a pre-embedded KG, narrowing down the candidate set for subsequent beam search. This approach addresses the latency and computational overhead typical of traditional agent-based methods. Path-RAG then constructs candidate reasoning steps by combining immediate semantic similarity with the structural connectivity of the graph, overcoming the dependence on highly precise retrievers in standard retrieval-based approaches. Next, the DVBS module employs an LLM-generated planning to guide the beam search that builds reasoning paths step-by-step. At each step, deductive verification checks whether the accumulated reasoning steps logically supports the user question, ensuring the final reasoning path is both verifiable and accurate.
@@ 3.1 Path-RAG: Reasoning Path Retrieval-Augmented Generation
Previous agent-based methods Ma et al. (2024); Sun et al. (2023) treat LLMs as agents that iteratively interact with KGs to find reasoning paths and answer, which necessitate multiple rounds of interaction between agents and KGs and lead to high computational costs and latency. We instead propose a module, Path-RAG, which iteratively pre-select a smaller candidate set to reduces the search space for exploring the potential reasoning paths from KGs. It consists of three steps and we detail the workflow as follows:
@@ 4.3 Robustness Analysis
To verify the faithfulness of our step-wise method, we conduct an error analysis regarding the whole reasoning path generation using RoG Luo et al. (2024). We quantify the validity of reasoning path using validity ratio (VR), which is defined as the ratio of reasoning steps that existed in the KG to the total number of the reasoning steps in the output reasoning path (i.e., $\mathrm{VR}=\frac{N_{\mathrm{valid-steps}}}{N_{\text{all-steps}}}$). As shown in Figure 3, only 67% of generated reasoning steps are valid, while the remaining 33% of reasoning steps either have a format error or do not exist in the KG. This illustrates that the reasoning steps generated offer few guarantees about feasibility especially when multiple consecutive steps are combined into a reasoning path. While our method leverage step-wise verification to ensure that each of the reasoning step exist in the KG and logically connected.
Figure 3: Analysis of reasoning errors in RoG Luo et al. (2024) over WebQSP.
@@ 5 Related Work
Large language models (LLMs) demonstrate impressive capabilities in reasoning tasks but often generate hallucinated or factually incorrect outputs, particularly in complex, multi-step scenarios Huang et al. (2025); Li et al. (2024); Chen et al. (2025); He et al. (2025). This unreliability reduces their effectiveness in knowledge-intensive applications. Knowledge graphs (KGs) have emerged as a solution by offering structured, verifiable data that supports transparent and multi-hop reasoning Sui et al. (2024). Unlike document-based retrieval-augmented generation approaches, KGs provide direct access to relational facts, enhancing both interpretability and traceability Chen et al. (2024).
@@
KG-enhanced reasoning methods are generally categorized into retrieval-based and agent-based models. Retrieval-based approaches, such as DeCAF Yu et al. (2023), rely on text-based retrieval to select relevant information from KGs and jointly generate answers and logical forms, but their performance can degrade without precise retrieval mechanisms. In contrast, agent-based models, like ToG Sun et al. (2023), iteratively explore reasoning paths but suffer from high computational overhead. To address these limitations, recent methods like RoG Luo et al. (2024) and GCR Luo et al. (2025) have sought to integrate KG structure into LLM training or decoding to improve reasoning fidelity and explanation generation. To improve the faithfulness of the LLM reasoning, KD-CoT Wang et al. (2023) verifies sub-reasoning steps through external KGs to prevent errors during inference, while NSM He et al. (2021) employs a teacher-student architecture to learn intermediate supervision signals that guide reasoning.
@@ A.1 Baselines
  • RoG Luo et al. (2024): incorporate the underling structure of KGs into LLMs throught pre-training and fine-tuning to generate the reasoning path and explanation.
  • GCR Luo et al. (2025) propose to integrate KG structure into the LLM decoding process to conduct graph-constrained reasoning.
@@ A.2 Datasets & Metrics
We consider three KGQA benchmark: WebQuestionSP (WebQSP) Yih et al. (2016), Complex WebQuestions (CWQ) Talmor and Berant (2018) and CR-LT-KGQA Guo et al. (2024) in this work. We follow previous work Luo et al. (2024) to use the same training and testing splits for fair comparison over WebQSP and CWQ. The questions from both WebQSP and CWQ can be reasoned using Freebase KGs [^2] https://github.com/microsoft/FastRDFStore. To address the bias in WebQSP and CWQ, which predominantly feature popular entities and there is a likelihood that their data might have been incorporated into the pre-training corpora of LLMs, we further test our method on CR-LT-KGQA (discussed in §A.2). We use the complete dataset from CR-LT-KGQA in our experiments, as it comprises only 200 samples. Each of the question can be reasoned based on the Wikidata [^3] https://www.wikidata.org/wiki/Wikidata:Main_Page. The statistics of the datasets are given in Table 10 and Table 9. To streamline the KGs, we follow RoG Luo et al. (2024) and utilize a subgraph of Freebase by extracting all triples that fall within the maximum reasoning hops from the question entities in WebQSP and CWQ. Similarly, we construct the corresponding sub-graphs of Wikidata for CR-LT-KGQA. We assess the performance of the methods by analyzing the F1 and Hits@1 metrics for CWQ and WebQSP, and by evaluating the accuracy for CR-LT-KGQA. The statistics of the datasets can be found in Table 9 and Table 10.
@@ A.3 Backbone LLMs & Embedding Methods
Embedding Methods. We assess the robustness of the retrieval module Path-RAG on different embedding models. We consider probabilistic ranking function like BM25444https://en.wikipedia.org/wiki/Okapi_BM25, dense retrieval using smaller language models like SentenceBERT [^5] https://sbert.net and E5666https://huggingface.co/intfloat/e5-large, and more advanced embedding model like text-embedding-3-small from OpenAI [^7] https://platform.openai.com/docs/guides/embeddings.